OpenStudy (anonymous):

find the antiderivative of cuberoot of x^2 + xsqrtx

hartnn (hartnn):

is this your question ? \(\Large \int \sqrt[3]{x^2 + x\sqrt x}dx\) ?? or

hartnn (hartnn):

\(\Large \int \sqrt[3]{x^2} + x\sqrt xdx\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the second one

hartnn (hartnn):

then its easy :) do you know \(\Large \int x^n dx =...?\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x^n+1/ n+1

hartnn (hartnn):

good so \(\Large \sqrt[3]{x^2} = x^{\dfrac{2}{3}}\) so n = 2/3

hartnn (hartnn):

\(\Large x \sqrt x = x^ {1+1/2} = x^{3/2}\) so for 2nd term n = 3/2

hartnn (hartnn):

got this ?

OpenStudy (anonymous):

yeah! thank you so much!

hartnn (hartnn):

welcome ^_^ and since you're new here \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)

OpenStudy (anonymous):

thanks!